# Model Reconstruction Using Counterfactual Explanations: Mitigating the Decision Boundary Shift

### v1

@@ Abstract
Counterfactual explanations find ways of achieving a favorable model outcome with minimum input perturbation. However, counterfactual explanations can also be exploited to steal the model by strategically training a surrogate model to give similar predictions as the original (target) model. In this work, we investigate model extraction by specifically leveraging the fact that the counterfactual explanations also lie quite close to the decision boundary. We propose a novel strategy for model extraction that we call Counterfactual Clamping Attack (CCA) which trains a surrogate model using a unique loss function that treats counterfactuals differently than ordinary instances. Our approach also alleviates the related problem of decision boundary shift that arises in existing model extraction attacks which treat counterfactuals as ordinary instances. We also derive novel mathematical relationships between the error in model approximation and the number of queries using polytope theory. Experimental results demonstrate that our strategy provides improved fidelity between the target and surrogate model predictions on several real world datasets.
@@ 1 Introduction
Counterfactual explanations (Wachter, Mittelstadt, and Russell 2017) are well-known for providing guidance on how to obtain a more favorable outcome from a machine learning model, e.g., increase your income by 10K to qualify for the loan. Given an input instance, a counterfactual explanation (also called a counterfactual) finds another instance that belongs to a different output class. Typically, the counterfactual is selected based on certain desirable criteria such as proximity to the original instance, limited feature changes, etc.
While counterfactuals provide several benefits, it has also been shown (Aïvodji, Bolot, and Gambs 2020) that counterfactuals have the potential to leak sensitive information about the underlying model, causing privacy concerns. It can become a serious threat, especially in Machine Learning as a Service (MLaaS)platforms (Gong et al. 2021; Tramèr et al. 2016) that are emerging as a popular alternative to learning models on site, allowing users to query the model for a specified cost (Gong et al. 2020; Juuti et al. 2019). Preserving the privacy of these models is a huge concern for the operators of such services because by strategically querying the model, an adversary may be able to “steal” the model (Wang, Qian, and Miao 2022; Gong et al. 2021; Tramèr et al. 2016; Pal et al. 2020; Juuti et al. 2019). Stealing usually involves training a surrogate model to provide similar predictions as the target model, a practice also referred to as model extraction. In this work, our main question is: How can counterfactuals be leveraged to perform model extraction attacks?
One existing approach is to treat counterfactuals as ordinary labeled points and use them for training the surrogate model (Aïvodji, Bolot, and Gambs 2020). While this may work for a well-balanced query dataset with queries from the two classes lying roughly equidistant to the decision boundary, it is not the same for unbalanced datasets. The surrogate decision boundary might not always overlap with that of the target model (see Fig. 1), a problem also referred to as a decision boundary shift. This is a result of the learning process where the boundary is kept as far as possible from the training examples (margin) to achieve better generalization (Shokri, Strobel, and Zick 2021). This issue is aggravated when the system provides only one-sided counterfactuals, i.e., counterfactuals only for queries with unfavorable predictions which is a common use case, e.g., only for rejected applicants in a loan application. This is because the counterfactuals are typically quite close to the decision boundary. Hence, when treated as ordinary labeled points for training the surrogate model, one-sided counterfactuals make the dataset unbalanced in terms of the distance from the decision boundary.
@@
Figure 1: Decision boundary shift in model extraction when counterfactuals are treated as ordinary labeled instances.
@@
In this work, we analyze how model extraction attacks can be refined by specifically leveraging the fact that the counterfactuals are quite close to the decision boundary. We demonstrate strategies that alleviate the decision-boundary-shift issue. We provide novel theoretical guarantees for the attacks, addressing an important knowledge gap in the existing literature. In contrast to existing attacks Aïvodji, Bolot, and Gambs (2020) and Wang, Qian, and Miao (2022) which require the system to provide counterfactuals for queries from both sides of the decision boundary, our methods require only one-sided counterfactuals (from one side of the decision boundary). Our contributions can be listed concisely as follows:
  • Fundamental guarantees on model extraction using closest counterfactuals: We first demonstrate how the closest counterfactuals can provide a linear approximation of the decision boundary around the counterfactuals (Lemma 3.1). Our key contribution lies in analyzing the goodness of this linear approximation using polytope theory. We provide a novel theoretical characterization of the relationship between the number of queries and the average agreement of the surrogate model (obtained using linear approximations) with the target model (see Theorem 3.2).
  • An attack based on the proximity of the counterfactuals to the target decision boundary: Noting the practical challenges associated with generating the closest counterfactual, we focus on an alternative strategy where we exploit only the fact that counterfactuals lie reasonably close to the decision boundary, but need not be exactly the closest. We first observe that the Lipschitz continuity of the target and surrogate models would constrict the deviations of their prediction probabilities around matching points (Theorem 3.4). To be precise, if $m_{1}$ and $m_{2}$ are two Lipschitz continuous classifiers (with Lipschitz constants $\gamma_{1}$ and $\gamma_{2}$, respectively) and $\boldsymbol{w}$ is a matching point, i.e., an input instance such that $m_{1}(\boldsymbol{w})=m_{2}(\boldsymbol{w})$, then for any $\boldsymbol{x}$ in the input space, $m_{1}(\boldsymbol{x})$ and $m_{2}(\boldsymbol{x})$ differ at most by $(\gamma_{1}+\gamma_{2})||\boldsymbol{x}-\boldsymbol{w}||_{2}$. This observation motivates a refined way of utilizing counterfactuals for model extraction. By forcing the surrogate model to match the prediction probabilities of the target model at a sufficiently large number of points along the decision boundary, we can force the decision boundaries of the target and surrogate models to overlap significantly. Furthermore, this can be precisely achieved with counterfactuals because they lie closer to the target decision boundary where the prediction probability of the target model is roughly 0.5. Accordingly, we propose a novel model extraction attack – that we call Counterfactual Clamping Attack (CCA) – which introduces a novel loss function for training the surrogate model by clamping its decision boundary to the target decision boundary at the counterfactuals (see Equation 3.2).
@@
  • Empirical validation: We conduct experiments on both synthetic datasets, as well as, four real-world datasets, namely, Adult Income (Becker and Kohavi 1996), COMPAS (Angwin et al. 2016), DCCC (Yeh 2016), and HELOC (FICO 2018). Our strategy outperforms the existing strategy in Aïvodji, Bolot, and Gambs (2020) over all these datasets (Section 4) using one-sided counterfactuals, i.e., counterfactuals only for queries from the unfavorable side of the decision boundary. We also include additional experiments to observe the effects of model architecture, Lipschitzness, and other types of counterfactual generation methods as well as an ablation study to experiment with other loss functions.
@@
Related Works: A plethora of counterfactual-generating mechanisms has been suggested in existing literature (Guidotti 2022; Barocas, Selbst, and Raghavan 2020; Verma et al. 2022; Karimi et al. 2022, 2020; Mothilal, Sharma, and Tan 2020; Dhurandhar et al. 2018; Deutch and Frost 2019; Mishra et al. 2021). Related works that focus on leaking information about the dataset from counterfactual explanations include membership inference attacks (Pawelczyk, Lakkaraju, and Neel 2023) and explanation-linkage attacks (Goethals, Sörensen, and Martens 2023). Shokri, Strobel, and Zick (2021) examines membership inference from other types of explanations, e.g., feature-based. Instead, we focus on model extraction from counterfactual explanations.
@@
Model extraction attacks (without counterfactuals) have been the topic of a wide array of studies (see surveys Gong et al. (2020) and Oliynyk, Mayer, and Rauber (2023)). Various mechanisms such as model inversion (Gong et al. 2021), equation solving (Tramèr et al. 2016), as well as active learning have been considered (Pal et al. 2020). Milli et al. (2019) looks into model reconstruction using other types of explanations, e.g., gradient-based. Yadav, Moshkovitz, and Chaudhuri (2023) explore algorithmic auditing using counterfactual explanations, focusing on linear classifiers and decision trees. Using counterfactual explanations for model extraction has received limited attention, with the notable exception of Aïvodji, Bolot, and Gambs (2020) and Wang, Qian, and Miao (2022). Aïvodji, Bolot, and Gambs (2020) suggest using counterfactuals as ordinary labeled examples while training the surrogate model. This suffers from the issue of decision boundary shift, particularly for unbalanced query datasets (one-sided counterfactuals). Wang, Qian, and Miao (2022) introduces a strategy of further querying for the counterfactual of the counterfactual. However, both these methods require the system to provide counterfactuals for queries from both sides of the decision boundary. Nevertheless, a user with a favorable decision may not usually require a counterfactual explanation, and hence a system providing one-sided counterfactuals might be more common, wherein lies our significance. While model extraction attacks (without counterfactuals) have received interest from a theoretical perspective (Tramèr et al. 2016; Papernot et al. 2017; Milli et al. 2019), model extraction attacks involving counterfactual explanations lack such a theoretical understanding. Our work proposes and theoretically analyzes model extraction attacks by explicitly utilizing the fact that the counterfactuals are close to the decision boundary. We provide novel performance guarantees in terms of query complexity for the proposed attacks, leveraging polytope theory, and also address the decision-boundary shift issue.
@@ 2 Preliminaries
Notations: We consider binary classification models $m$ that take an input value $\boldsymbol{x}\in\mathbb{R}^{d}$ and output a probability between 00 and $1$. The final predicted class is denoted by $\left\lfloor{m(\boldsymbol{x})}\right\rceil\in\{0,1\}$ which is obtained by thresholding the output probability at $0.5$ as follows: $\left\lfloor{m(\boldsymbol{x})}\right\rceil=\mathds{1}[m(\boldsymbol{x})\geq 0.5]$ where $\mathds{1}[\cdot]$ denotes the indicator function. Throughout the paper, we denote the output probability by $m(\boldsymbol{x})$ and the corresponding thresholded output by $\left\lfloor{m(\boldsymbol{x})}\right\rceil$. Consequently, the decision boundary of the model $m$ is the $(d-1)$-dimensional hypersurface (generalization of surface in higher dimensions; see Definition 2.4) in the input space, given by $\partial{\mathbb{M}}=\{\boldsymbol{x}:m(\boldsymbol{x})=0.5\}$. We call the region where $\left\lfloor{m(\boldsymbol{x})}\right\rceil=1$ as the favorable region and the region where $\left\lfloor{m(\boldsymbol{x})}\right\rceil=0$ as the unfavorable region. We always state the convexity/concavity of the decision boundary with respect to the favorable region (i.e., the decision boundary is convex if the set ${\mathbb{M}}=\{\boldsymbol{x}\in{\mathbb{R}}^{d}:\left\lfloor{m(\boldsymbol{x})}\right\rceil=1\}$ is convex). We assume that upon knowing the range of values for each feature, the $d-$dimensional input space can be normalized so that the inputs lie within the set $[0,1]^{d}$ (the $d-$dimensional unit hypercube), as is common in literature (Liu et al. 2020; Tramèr et al. 2016; Hamman et al. 2023; Black, Wang, and Fredrikson 2022). We let $g_{m}$ denote the counterfactual generation mechanism corresponding to the model $m$.
@@
Given a cost function $c:[0,1]^{d}\times[0,1]^{d}\rightarrow{\mathbb{R}}^{+}_{0}$ for measuring the quality of a counterfactual, and a model $m$, the corresponding counterfactual generating mechanism is the mapping $g_{m}:[0,1]^{d}\rightarrow[0,1]^{d}$ specified as
$g_{m}(\boldsymbol{x})=\operatorname*{\arg\;\min}_{\begin{subarray}{c}\boldsymbol{w}\in[0,1]^{d},\ \left\lfloor{m(\boldsymbol{x})}\right\rceil\neq\left\lfloor{m(\boldsymbol{w})}\right\rceil\end{subarray}}c(\boldsymbol{x},\boldsymbol{w}).$ (1)
The cost $c(\boldsymbol{x},\boldsymbol{w})$ is selected based on specific desirable criteria, e.g., $c(\boldsymbol{x},\boldsymbol{w})=||\boldsymbol{x}-\boldsymbol{w}||_{p}$, with $||\cdot||_{p}$ denoting the $L_{p}$-norm. Specifically, $p=2$ leads to the following definition of the closest counterfactual (Wachter, Mittelstadt, and Russell 2017; Laugel et al. 2017; Mothilal, Sharma, and Tan 2020).
@@
Problem Setting: Our problem setting involves a target model $m$ which is pre-trained and assumed to be hosted on a MLaaS platform (see Fig. 2). Users can query it with a set of
@@
input instances ${\mathbb{D}}\subseteq[0,1]^{d}$ through an Application Programming Interface (API). The API will provide the users with the set of predictions, i.e., $\{\left\lfloor{m(\boldsymbol{x})}\right\rceil:\boldsymbol{x}\in{\mathbb{D}}\}$, and a set of one-sided counterfactuals for the instances whose predicted class is 00, i.e., $\{g_{m}(\boldsymbol{x}):\boldsymbol{x}\in{\mathbb{D}},\left\lfloor{m(\boldsymbol{x})}\right\rceil=0\}$. Note that, by the definition of a counterfactual, $\left\lfloor{m(g_{m}(\boldsymbol{x}))}\right\rceil=1$ for all $\boldsymbol{x}$ with $\left\lfloor{m(\boldsymbol{x})}\right\rceil=0$. An adversary, while appearing to be a normal user, can query with an attack dataset ${\mathbb{D}}_{\text{attack}}\subseteq[0,1]^{d}$ and use the returned labels and counterfactuals to train their own surrogate model $\tilde{m}$. The goal of the adversary is to achieve a certain level of performance with as few queries as possible. In this work, we use fidelity as our performance metric for model extraction [^1] Performance can be evaluated using accuracy or fidelity (Jagielski et al. 2020). Accuracy is a measure of how well the surrogate model can predict the true labels over the data manifold of interest. Fidelity measures the agreement between labels predicted by the surrogate and the target models. While attacks based on both measures have been proposed in literature, fidelity-based attacks have been deemed more useful as a first step in designing future attacks (Jagielski et al. 2020; Papernot et al. 2017)..
Definition 2.3 (Fidelity (Aïvodji, Bolot, and Gambs 2020)).
@@
Next, we show that touching hypersurfaces share a common tangent hyperplane at their point of contact. This result is instrumental in exploiting the closest counterfactuals in a model extraction attack (proof in Appendix A.1).
@@
Lipschitz Continuity: Lipschitz continuity, a property that is often encountered in related works (Bartlett, Foster, and Telgarsky 2017; Gouk et al. 2021; Pauli et al. 2021; Hamman et al. 2023; Liu et al. 2020; Marques-Silva et al. 2021), can be exploited in perfecting model extraction attacks.
Definition 2.7 (Lipschitz Continuity).
A model $m$ is said to be Lipschitz continuous with a constant $\gamma$ if
$|m(\boldsymbol{x}_{1})-m(\boldsymbol{x}_{2})|\leq\gamma||\boldsymbol{x}_{1}-\boldsymbol{x}_{2}||_{2}$ (2)
for all $\boldsymbol{x}_{1},\boldsymbol{x}_{2}\in[0,1]^{d}$, with $\gamma\in{\mathbb{R}}_{0}^{+}$.

### v2

# Model Reconstruction Using Counterfactual Explanations: A Perspective From Polytope Theory

@@ Abstract
Counterfactual explanations provide ways of achieving a favorable model outcome with minimum input perturbation. However, counterfactual explanations can also be leveraged to reconstruct the model by strategically training a surrogate model to give similar predictions as the original (target) model. In this work, we analyze how model reconstruction using counterfactuals can be improved by further leveraging the fact that the counterfactuals also lie quite close to the decision boundary. Our main contribution is to derive novel theoretical relationships between the error in model reconstruction and the number of counterfactual queries required using polytope theory. Our theoretical analysis leads us to propose a strategy for model reconstruction that we call Counterfactual Clamping Attack (CCA) which trains a surrogate model using a unique loss function that treats counterfactuals differently than ordinary instances. Our approach also alleviates the related problem of decision boundary shift that arises in existing model reconstruction approaches when counterfactuals are treated as ordinary instances. Experimental results demonstrate that our strategy improves fidelity between the target and surrogate model predictions on several datasets.
@@ 1 Introduction
Counterfactual explanations (also called counterfactuals) have emerged as a burgeoning area of research (Wachter et al., 2017; Guidotti, 2022; Barocas et al., 2020; Verma et al., 2022; Karimi et al., 2022) for providing guidance on how to obtain a more favorable outcome from a machine learning model, e.g., increase your income by 10K to qualify for the loan. Interestingly, counterfactuals can also reveal information about the underlying model, posing a nuanced interplay between model privacy and explainability (Aïvodji et al., 2020; Wang et al., 2022). Our work provides the first theoretical analysis between the error in model reconstruction using counterfactuals and the number of counterfactuals queried for, through the lens of polytope theory.
Model reconstruction using counterfactuals can have serious implications in Machine Learning as a Service (MLaaS)platforms that allow users to query a model for a specified cost (Gong et al., 2020). An adversary may be able to “steal” the model by querying for counterfactuals and training a surrogate model to provide similar predictions as the target model, a practice also referred to as model extraction. On the other hand, model reconstruction could also be beneficial for preserving applicant privacy, e.g., if an applicant wishes to evaluate their chances of acceptance from crowdsourced information before formally sharing their own application information with an institution, often due to resource constraints or having a limited number of attempts to apply (e.g., applying for credit cards reduces the credit score (Capital One, 2024)). Our goal is to formalize how faithfully can one reconstruct an underlying model given a set of counterfactual queries.
@@
Figure 1: Decision boundary shift when counterfactuals are treated as ordinary labeled points.
An existing approach for model reconstruction is to treat counterfactuals as ordinary labeled points and use them for training a surrogate model (Aïvodji et al., 2020). While this may work for a well-balanced counterfactual queries from the two classes lying roughly equidistant to the decision boundary, it is not the same for unbalanced datasets. The surrogate decision boundary might not always overlap with that of the target model (see Fig. 1), a problem also referred to as a decision boundary shift. This is due to the learning process where the boundary is kept far from the training examples (margin) for better generalization (Shokri et al., 2021).
The decision boundary shift is aggravated when the system provides only one-sided counterfactuals, i.e., counterfactuals only for queries with unfavorable predictions. If one were allowed to query for two-sided counterfactuals, the decision boundary shift may be tackled by querying for the counterfactual of the counterfactual (Wang et al., 2022). However, such strategies cannot be applied when only one-sided counterfactuals are available, which is more common and also a more challenging use case for model reconstruction, e.g., counterfactuals are only available for the rejected applicants to get accepted for a loan but not the other way.
@@
In this work, we analyze how model reconstruction using counterfactuals can be improved by leveraging the fact that the counterfactuals are quite close to the decision boundary. We provide novel theoretical analysis for model reconstruction using polytope theory, addressing an important knowledge gap in the existing literature. We demonstrate reconstruction strategies that alleviate the decision-boundary-shift issue for one-sided counterfactuals. In contrast to existing strategies Aïvodji et al. (2020) and Wang et al. (2022) which require the system to provide counterfactuals for queries from both sides of the decision boundary, we are able to reconstruct using only one-sided counterfactuals, a problem that we also demonstrate to be theoretically more challenging than the two-sided case (see Corollary 3.8). In summary, our contributions can be listed as follows:
  • Fundamental guarantees on model reconstruction using counterfactuals: We derive novel theoretical relationships between the error in model reconstruction and the number of counterfactual queries (query complexity) under three settings: (i) Convex decision boundaries and closest counterfactuals (Theorem 3.2): We rely on convex polytope approximations to derive an exact relationship between expected model reconstruction error and query complexity; (ii) ReLU networks and closest counterfactuals (Theorem 3.6): Relaxing the convexity assumption, we provide probabilistic guarantees on the success of model reconstruction as a function of number of counterfactual queries; and (iii) Beyond closest counterfactuals: We provide approximate guarantees for a broader class of models, including ReLU networks and locally-Lipschitz continuous models (Theorem 3.10).
  • Model reconstruction strategy with unique loss function: We devise a reconstruction strategy – that we call Counterfactual Clamping Attack (CCA) – that exploits only the fact that the counterfactuals lie reasonably close to the decision boundary, but need not be exactly the closest.
@@
  • Empirical validation: We conduct experiments on both synthetic datasets, as well as, four real-world datasets, namely, Adult Income (Becker and Kohavi, 1996), COMPAS (Angwin et al., 2016), DCCC (Yeh, 2016), and HELOC (FICO, 2018). Our strategy outperforms the existing baseline (Aïvodji et al., 2020) over all these datasets (Section 4) using one-sided counterfactuals, i.e., counterfactuals only for queries from the unfavorable side of the decision boundary. We also include additional experiments to observe the effects of model architecture, Lipschitzness, and other types of counterfactual generation methods as well as ablation studies with other loss functions. A python-based implementation is available at: https://github.com/pasandissanayake/model-reconstruction-using-counterfactuals.
@@
Related Works: A plethora of counterfactual-generating mechanisms has been suggested in existing literature (Guidotti, 2022; Barocas et al., 2020; Verma et al., 2022; Karimi et al., 2022, 2020; Mothilal et al., 2020; Dhurandhar et al., 2018; Deutch and Frost, 2019; Mishra et al., 2021). Related works that focus on leaking information about the dataset from counterfactual explanations include membership inference attacks (Pawelczyk et al., 2023) and explanation-linkage attacks (Goethals et al., 2023). Shokri et al. (2021) examine membership inference from other types of explanations, e.g., feature-based. Model reconstruction (without counterfactuals) has been the topic of a wide array of studies (see surveys Gong et al. (2020) and Oliynyk et al. (2023)). Various mechanisms such as equation solving (Tramèr et al., 2016) and active learning have been considered (Pal et al., 2020).
Model inversion (Gong et al., 2021; Struppek et al., 2022; Zhao et al., 2021) is another form of extracting information about a black box model, under limited access to the model aspects. In contrast to model extraction where the goal is to replicate the model itself, in model inversion an adversary tries to extract the representative attributes of a certain class with respect to the target model. In this regard, Zhao et al. (2021) focuses on exploiting explanations for image classifiers such as saliency maps to improve model inversion attacks. Struppek et al. (2022) proposes various methods based on Generative Adversarial Networks to make model inversion attacks robust (for instance, to distributional shifts) in the domain of image classification.
@@
Milli et al. (2019) looks into model reconstruction using other types of explanations, e.g., gradients. Yadav et al. (2023) explore algorithmic auditing using counterfactual explanations, focusing on linear classifiers and decision trees. Using counterfactual explanations for model reconstruction has received limited attention, with the notable exceptions of Aïvodji et al. (2020) and Wang et al. (2022). Aïvodji et al. (2020) suggest using counterfactuals as ordinary labeled examples while training the surrogate model, leading to decision boundary shifts, particularly for unbalanced query datasets (one-sided counterfactuals). Wang et al. (2022) introduces a strategy of mitigating this issue by further querying for the counterfactual of the counterfactual. However, both these methods require the system to provide counterfactuals for queries from both sides of the decision boundary. Nevertheless, a user with a favorable decision may not usually require a counterfactual explanation, and hence a system providing one-sided counterfactuals might be more common, wherein lies our significance. While model reconstruction (without counterfactuals) has received interest from a theoretical perspective (Tramèr et al., 2016; Papernot et al., 2017; Milli et al., 2019), model reconstruction involving counterfactual explanations lack such a theoretical understanding. Our work theoretically analyzes model reconstruction using polytope theory and proposes novel strategies thereof, also addressing the decision-boundary shift issue.
@@ 2 Preliminaries
Notations: We consider binary classification models $m$ that take an input value $\boldsymbol{x}\in\mathbb{R}^{d}$ and output a probability between 00 and $1$. The final predicted class is denoted by $\left\lfloor{m(\boldsymbol{x})}\right\rceil\in\{0,1\}$ which is obtained by thresholding the output probability at $0.5$ as follows: $\left\lfloor{m(\boldsymbol{x})}\right\rceil=\mathds{1}[m(\boldsymbol{x})\geq 0.5]$ where $\mathds{1}[\cdot]$ denotes the indicator function. Throughout the paper, we denote the output probability by $m(\boldsymbol{x})$ and the corresponding thresholded output by $\left\lfloor{m(\boldsymbol{x})}\right\rceil$. Consequently, the decision boundary of the model $m$ is the $(d-1)$-dimensional hypersurface (generalization of surface in higher dimensions; see Definition 2.5) in the input space, given by $\partial{\mathbb{M}}=\{\boldsymbol{x}:m(\boldsymbol{x})=0.5\}$. We call the region where $\left\lfloor{m(\boldsymbol{x})}\right\rceil=1$ as the favorable region and the region where $\left\lfloor{m(\boldsymbol{x})}\right\rceil=0$ as the unfavorable region. We always state the convexity/concavity of the decision boundary with respect to the favorable region (i.e., the decision boundary is convex if the set ${\mathbb{M}}=\{\boldsymbol{x}\in{\mathbb{R}}^{d}:\left\lfloor{m(\boldsymbol{x})}\right\rceil=1\}$ is convex). We assume that upon knowing the range of values for each feature, the $d-$dimensional input space can be normalized so that the inputs lie within the set $[0,1]^{d}$ (the $d-$dimensional unit hypercube), as is common in literature (Liu et al., 2020; Tramèr et al., 2016; Hamman et al., 2023; Black et al., 2022). We let $g_{m}$ denote the counterfactual generating mechanism corresponding to the model $m$.
@@
Given a cost function $c:[0,1]^{d}\times[0,1]^{d}\rightarrow{\mathbb{R}}^{+}_{0}$ for measuring the quality of a counterfactual, and a model $m$, the corresponding counterfactual generating mechanism is the mapping $g_{m}:[0,1]^{d}\rightarrow[0,1]^{d}$ specified as follows: $g_{m}(\boldsymbol{x})=\operatorname*{\arg\;\min}_{\boldsymbol{w}\in[0,1]^{d}}c(\boldsymbol{x},\boldsymbol{w}),$ such that $\left\lfloor{m(\boldsymbol{x})}\right\rceil\neq\left\lfloor{m(\boldsymbol{w})}\right\rceil$.
The cost $c(\boldsymbol{x},\boldsymbol{w})$ is selected based on specific desirable criteria, e.g., $c(\boldsymbol{x},\boldsymbol{w})=||\boldsymbol{x}-\boldsymbol{w}||_{p}$, with $||\cdot||_{p}$ denoting the $L_{p}$-norm. Specifically, $p=2$ leads to the following definition of the closest counterfactual (Wachter et al., 2017; Laugel et al., 2017; Mothilal et al., 2020).
@@
Given a model $m$ and a counterfactual generating method $g_{m}$, we define the inverse counterfactual region ${\mathbb{G}}$ for a subset ${\mathbb{H}}\subseteq[0,1]^{d}$ to be the region whose counterfactuals under $g_{m}$ fall in ${\mathbb{H}}$.
Definition 2.3 (Inverse Counterfactual Region).
The inverse counterfactual region ${\mathbb{G}}_{m,g_{m}}$ of ${\mathbb{H}}\subseteq[0,1]^{d}$ is the the region defined as: ${\mathbb{G}}_{m,g_{m}}({\mathbb{H}})=\{\boldsymbol{x}\in[0,1]^{d}:g_{m}(\boldsymbol{x})\in{\mathbb{H}}\}.$
@@
Problem Setting: Our problem setting involves a target model $m$ which is pre-trained and assumed to be hosted on a MLaaS platform (see Fig. 2). Any user can query it with a set of input instances ${\mathbb{D}}\subseteq[0,1]^{d}$ (also called counterfactual queries) and will be provided with the set of predictions, i.e., $\{\left\lfloor{m(\boldsymbol{x})}\right\rceil:\boldsymbol{x}\in{\mathbb{D}}\}$, and a set of one-sided counterfactuals for the instances whose predicted class is 00, i.e., $\{g_{m}(\boldsymbol{x}):\boldsymbol{x}\in{\mathbb{D}},\left\lfloor{m(\boldsymbol{x})}\right\rceil=0\}$. Note that, by the definition of a counterfactual, $\left\lfloor{m(g_{m}(\boldsymbol{x}))}\right\rceil=1$ for all $\boldsymbol{x}$ with $\left\lfloor{m(\boldsymbol{x})}\right\rceil=0$. The goal of the user is to train a surrogate model to achieve a certain level of performance with as few queries as possible. In this work, we use fidelity as our performance metric for model reconstruction [^1] Performance can be evaluated using accuracy or fidelity (Jagielski et al., 2020). Accuracy is a measure of how well the surrogate model can predict the true labels over the data manifold of interest. While attacks based on both measures have been proposed in literature, fidelity-based attacks have been deemed more useful as a first step in designing future attacks (Jagielski et al., 2020; Papernot et al., 2017)..
Definition 2.4 (Fidelity (Aïvodji et al., 2020)).
@@
Next, we show that touching hypersurfaces share a common tangent hyperplane at their point of contact. This result is instrumental in exploiting the closest counterfactuals in model reconstruction (proof in Appendix A.1).